Solve by elimination method 2r - 5s = -14 5r + 2s = 52 what is the solution of the system
Answer by Maths68(1474)

2r - 5s = -14......(1)
5r + 2s = 52.......(2)
Multiply (1) by 5 and (2) by 2
10r-25s=-70.....(3)
10r+4s=104.......(4)
Subtract (4) from (3)
10r-25s=-70
-10r-4s=-104
------------------
-29s=-174
-29s/-29=-174/-29
s=6
Put the value of s in (1)
2r - 5s = -14......(1)
2r - 5(6) = -14
2r - 30 = -14
2r=-14+30
2r=16
2r/2=16/2
r=8
Solution Set
(r,s)=(8,6)




Check
=====
Put the value of r and s in (2)
5r + 2s = 52.......(2)
5(8) + 2(6) = 52
40+12=58
52=52